Lou Barlow, a divisional manager for Sage Company, has an opportunity to manufacture and sell one of two new products for a five-year period. His annual pay raises are determined by his division’s return on investment (ROI), which has exceeded 23% each of the last three years. He has computed the cost and revenue estimates for each product as follows:

Product A

Product B

Initial investment:

Cost of equipment (zero salvage value)

$

290,000

$

490,000

Annual revenues and costs:

Sales revenues

$

340,000

$

440,000

Variable expenses

$

154,000

$

206,000

Depreciation expense

$

58,000

$

98,000

Fixed out-of-pocket operating costs

$

79,000

$

59,000

The company’s discount rate is 15%.

Required:

1. Calculate the payback period for each product.

2. Calculate the net present value for each product.

3. Calculate the internal rate of return for each product.

4. Calculate the project profitability index for each product.

5. Calculate the simple rate of return for each product.

6a. For each measure, identify whether Product A or Product B is preferred.

6b. Based on the simple rate of return, Lou Barlow would likely:
Answer #1

Year А 3 Calculation of internal rate of return Product A Present value@ 15% в C=A/(1+15%) A 0 (290,000) 11 107,000 21 107,00Product B 2 Calculation of net present value of each product Product A Present value@ 15% C=A/(1+15%)^A (290,000) 107,000 1071. Calculation of pay back period Year Product A Present value@ 15% Cumulative present value C=A/(1+15%)^A 0 (290,000) (290,0